# A quasilinear Keller-Segel model with saturated discontinuous advection

### v2

@@ 4.4. Obtaining the initial data $\rho_{{\rm in}}$ for $\rho$
[…]
for a.e. $t_{1},t_{2}\in[0,T]$. It follows that the function $t\mapsto\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}^{p}$ is in $W^{1,1}(0,T)$, hence up to changing the values of $\rho:[0,T]\to L^{p}(\mathbb{R}^{d})$ in a zero measure set it holds that $t\mapsto\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}^{p}$ is continuous in $[0,T]$. As a consequence the bound $\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}$ holds for every $t\in[0,T]$. In particular $\rho(t)$ is weakly convergent (up to subsequences) in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$; however, its weak limit has to be $\rho_{0}$ due to (44), hence $\rho(t)\rightharpoonup\rho_{0}$ weakly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$ (no longer up to subsequences). It follows
$\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}\leq\liminf_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\limsup_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})},$
implying that $\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\to\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}$ as $t\to 0$. This fact and the weak convergence $\rho(t)\rightharpoonup\rho_{0}$ in $L^{p}(\mathbb{R}^{d})$, imply that $\rho(t)\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$.
4.5. Uniqueness
Let $\rho_{1}$ and $\rho_{2}$ be two solutions to (10) with the same initial data. As in Section 4, we subtract equation (10) for $\rho_{2}$ tested with $\varphi_{2}=\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}{-1}=\frac{\rho_{1}-\rho_{2}}{\rho_{2}+\sigma}$ to the equation for $\rho_{1}$ tested with $\varphi_{1}=\log\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)$. Following the same computations as in (4.1), we get
$=\int_{0}^{T}\frac{d}{dt}\int_{\mathbb{R}^{d}}(\rho_{2}+\sigma)\left[\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\log{\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)}-\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)+1\right]\,{\rm d}x\,{\rm d}t$
$=\int_{\mathbb{R}^{d}}(\rho_{2}+\sigma)\left[\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\log{\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)}-\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)+1\right](T)\,{\rm d}x.$
After some manipulations, this yields
$\int_{\mathbb{R}^{d}}(\rho_{2}+\sigma)\left[\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\log{\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)}-\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)+1\right](T)\,{\rm d}x$ (45)
$=-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{1}+\sigma)\left[\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\cdot$
@@ 4.5. Uniqueness
$\;\;\;-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}(\rho_{1}+\sigma)\left|\nabla\log\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
$\;\;\;+\xi_{1}+\xi_{2}$
${\leq\xi_{1}+\xi_{2}},$
@@
$\xi_{1}:=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\nabla\rho_{1}\nabla\left(\log\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\,{\rm d}x\,{\rm d}t,$
@@
$\xi_{2}:=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\nabla\rho_{2}\nabla\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\,{\rm d}x\,{\rm d}t.$
We want to show that $\limsup_{\sigma\to 0}(\xi_{1}+\xi_{2})\leq 0$. We estimate first $\xi_{1}$ and then $\xi_{2}$.
$\xi_{1}\leq$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\nabla\rho_{1}\frac{\nabla\rho_{2}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t$
$\leq$ $\underbrace{\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{1}|^{2}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t}_{\text{(I)}}$
$+\underbrace{\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{2}|^{2}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t}_{\text{(II)}}.$
@@
[…]
@@
$\xi_{2}\leq$ $\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{\nabla\rho_{2}\nabla\rho_{1}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{(\rho_{1}+\sigma)|\nabla\rho_{2}|^{2}}{(\rho_{2}+\sigma)^{2}}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{1}|^{2}}{\rho_{1}+\sigma}\,{\rm d}x\,{\rm d}t$
$-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{(\rho_{1}+\sigma)|\nabla\rho_{2}|^{2}}{(\rho_{2}+\sigma)^{2}}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;{\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{1}|^{2}}{\rho_{1}}\,{\rm d}x\,{\rm d}t}$
@@
This implies that $\rho_{1}=\rho_{2}$.
5. Disclosure statement
There are no relevant financial or non-financial competing interests to report.

### v3

@@ 4.5. Uniqueness
$\;\;\;-\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}(\rho_{1}+\sigma)\left|\nabla\log\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\right|^{2}\left[\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}+\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\,{\rm d}x\,{\rm d}t$
$\;\;\;+\xi_{1}+\xi_{2},$
@@
$\xi_{1}:=-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\nabla\rho_{1}\nabla\left(\log\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\,{\rm d}x\,{\rm d}t,$
@@
$\xi_{2}:=\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\nabla\rho_{2}\nabla\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\,{\rm d}x\,{\rm d}t.$
The first two integrals in (54) are positive, the first because of monotonicity and the second because it is just the product of positive functions. Since both have negative signs in front, we can ignore them and we get
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\left[\varphi_{1}\partial_{t}\rho_{1}-\varphi_{2}\partial_{t}\rho_{2}\right]\,{\rm d}x\;dt\leq\xi_{1}+\xi_{2}.$
We want to show that $\limsup_{\sigma\to 0}(\xi_{1}+\xi_{2})\leq 0$. Dominated convergence theorem will be used. We only have to show that the integrands in $\xi_{1}$ and $\xi_{2}$ are dominated by an integrable function (uniformly in $\sigma$). We start with $\xi_{1}$:
$\xi_{1}\leq$ $\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\nabla\rho_{1}\frac{\nabla\rho_{2}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t$
$\leq$ $\underbrace{\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{1}|^{2}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t}_{\text{(I)}}$
$+\underbrace{\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{1}}{\lvert\nabla\rho_{1}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{1}+\sigma)}{\lvert\nabla\rho_{1}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{2}|^{2}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t}_{\text{(II)}}.$
@@
$|\mbox{Integrand in (I)}|\leq\frac{1}{2}|\nabla\rho_{1}|\leq|\nabla\sqrt{\rho}_{1}|\sqrt{\rho_{1}}\in L^{1}(\mathbb{R}^{d}\times(0,T)).$
@@
$\xi_{2}\leq$ $\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{\nabla\rho_{2}\nabla\rho_{1}}{\rho_{2}+\sigma}\,{\rm d}x\,{\rm d}t$
$-\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{(\rho_{1}+\sigma)|\nabla\rho_{2}|^{2}}{(\rho_{2}+\sigma)^{2}}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{1}|^{2}}{\rho_{1}+\sigma}\,{\rm d}x\,{\rm d}t$
$-\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{(\rho_{1}+\sigma)|\nabla\rho_{2}|^{2}}{(\rho_{2}+\sigma)^{2}}\,{\rm d}x\,{\rm d}t$
$\leq$ $\;{\frac{1}{2}\int_{0}^{T}\hskip-5.69054pt\int_{\mathbb{R}^{d}}\left[\left(1-\frac{\chi\rho_{2}}{\lvert\nabla\rho_{2}\rvert}\right)_{+}-\left(1-\frac{\chi(\rho_{2}+\sigma)}{\lvert\nabla\rho_{2}\rvert}\right)_{+}\right]\frac{|\nabla\rho_{1}|^{2}}{\rho_{1}}\,{\rm d}x\,{\rm d}t.}$
@@
This implies that $\rho_{1}=\rho_{2}$ and finishes the proof of Theorem 1.